OpenStudy (anonymous):

Which expression is equivalent to four? A. (10 – 4) + 2 B. 10 – 4 + 2 C. 10 – (4 + 2) D.(10 – 4 + 2)

OpenStudy (nathan917):

C. 10 - (4+2) (4+2) = 6-10= 4

OpenStudy (uri):

@Young_Hitta Try solving it yourself first.

OpenStudy (nathan917):

I Agree, This was very easy.

OpenStudy (nathan917):

Anything with ( ) around it is first.

OpenStudy (uri):

@nathan917 Don't just give out the answers. Ask the user to do it themselves,specially for these. Thanks.

OpenStudy (nathan917):

@uri Will do, i would have explained it but there was nothing to explain.